# Thermal conductivity minimum with thickness in ultrathin films

### v2

@@ Abstract
In future work, this theory can be extended in several directions such as the inclusion of phonon transmission factors between the thin film and the substrates, by systematic studying the effect of the various parameters in the perspective of developing effective phonon engineering strategies for nanomaterials. Also, it will be interesting to apply the theory to more exotic materials such as few-layer 2D van der Waals materials, graphene, BN, MoS2, and amorphous and crystalline membranes. Important applications of the proposed framework to heat management of quantum computing systems are also envisaged.

### v3

@@ VIII Conclusions
Instead, at larger thickness values than the point of minimum, the thermal conductivity increases monotonically with the thickness $L$ simply because more phonon modes are available to carry heat (as the two hole-pocket spheres in the left panel of Fig. 1 keep shrinking with increasing $L$). This is consistent with the linear increase of the specific heat with $L$, $C_{v}(\omega)\sim L$, predicted by the phonon-confinement theory in Zaccone (2024).
In future work, this theory can be extended in several directions such as the inclusion of phonon transmission factors between the thin film and the substrates, by systematic studying the effect of the various parameters in the perspective of developing effective phonon engineering strategies for nanomaterials. The present framework can also be applied to experimental data in the thicker range $5<L<500$ nm, as reported in Cuffe et al. (2015); Chávez-Angel et al. (2014); Vega-Flick et al. (2016), in future work. Also, it will be interesting to apply the theory to more exotic materials such as few-layer 2D van der Waals materials, graphene, BN, MoS2, and amorphous and crystalline membranes, e.g. by including flexural and Lamb modes Kühn and Maasilta (2006) and disorder effects in the DOS Baggioli and Zaccone (2020). Important applications of the proposed framework to heat management of quantum computing systems are also envisaged.
Author Declarations
The authors have no conflicts to disclose.
Data Availability
Data sharing is not applicable to this article as no new data were created or analyzed in this study.